OpenStudy (gracejaylene(1)):

Solve the following automotive-services problem. In first gear, or low gear, an automobile's engine runs about three times as fast as the drive shaft. In second gear, the engine does not have to run as fast; usually it runs about 1.6 times faster than the drive shaft. Finally, in third, or high gear, the engine runs at the same speed as the drive shaft. Engine speed = 3,500 r.p.m. Transmission in first gear Drive-shaft speed (to nearest r.p.m.) =.

OpenStudy (gracejaylene(1)):

@Qwertty123 @-SpencerBraz-

OpenStudy (qwertty123):

I am not sure :(

OpenStudy (gracejaylene(1)):

@-SpencerBraz- HALP

OpenStudy (gracejaylene(1)):

@SkyVoltage

OpenStudy (phi):

an automobile's engine runs about three times as fast as the drive shaft. engine rpm = 3 * shaft rpm

OpenStudy (phi):

replace "engine rpm" with 3500 rpm 3* shaft rpm = 3500 rpm divide both sides by 3 3/3 * shaft rpm = 3500/3 rpm or shaft rpm = 3500/3 rpm